OpenStudy (anonymous):

sqare root (-252/4) WHY IS THE ANSWER 3i sort (7) IMAGINARY NUMBERS HELP PLEASE

OpenStudy (anonymous):

You cannot square root a negative number without resorting to imaginary numbers. -252/4 is equal to -63, and when you square root it, is 3i sqrt 7. 3 root 7 is equal to root 63.

OpenStudy (kinggeorge):

\[\sqrt{-252/4}=\sqrt{-63}=\sqrt{9\cdot(-7)}=3\sqrt{-7}\]\[3\sqrt{-7}=3\sqrt{-1}\sqrt{7}=3i \sqrt7\]

OpenStudy (kinggeorge):

You can factor the -1 out of the expression, and write it as \(i\) to get your solution.

OpenStudy (anonymous):